# Out of spuriousity: Improving robustness to spurious correlations without group annotations

### v1

@@ A.4 Feature Visualization
Figure 6: Embedding space of ERM (first row), DCWP (second row), and our pruned network (third row) on CelebA for predicting blond hair color. We visualize the train set embedding space with colors representing the class of non-blond hair (left column) and three other attributes: female, young, and beard (center column). To show how the approach works on test set, we visualize the class non-blond hair of test set, coloring above spurious attributes (right column). In both train and test embedding, ERM clusters well spurious attributes even those attribute labels are not used during training; showing its priority to learn spurious features. DCWP leveraging the ERM-misclassified cases can not mitigate the influence of attribute female, shown by a high purity cluster of samples with attribute female in both train and test representation space. Our pruned model mixes up samples in the same class, eliminating the biased learning behavior.

### v2

@@ A.4 Feature Visualization
Figure 6: Embedding space of ERM (first row), DCWP (second row), and PruSC (third row) on CelebA for predicting blond hair color. We visualize the train set embedding space with colors representing the class of non-blond hair (left column) and three other attributes: female, young, and beard (center column). To show how the approach works on test set, we visualize the class non-blond hair of test set, coloring above spurious attributes (right column). In both train and test embedding, ERM clusters well spurious attributes even those attribute labels are not used during training; showing its priority to learn spurious features. DCWP leveraging the ERM-misclassified cases can not mitigate the influence of attribute female, shown by a high purity cluster of samples with attribute female in both train and test representation space. Our pruned model mixes up samples in the same class, eliminating the biased learning behavior.
A.5 PruSC with (pseudo) Group Labels
In this section, we study the effectiveness of PruSC with annotated spurious attributes and other pseudo group labeling approaches such as using minority and majority groups based on ERM predictions Nam et al. (2020); Park et al. (2023); Zhang et al. (2022).
We conduct experiments on the CelebA dataset using the same frozen ERM model $f$ trained as defined in Sec. 5.3. The hyper-parameters remain identical to those used for training the original PruSC. Regarding the changes, first, we do not perform representation clustering and sub-sampling. Instead, we design $\mathcal{D}_{\text{task}}$ to be (pseudo) group-balanced and fine-tune the model after pruning with this $\mathcal{D}_{\text{task}}$. Second, the contrastive loss defined during the binary mask training stage remains the same, but with the new clusters definition, as described in the following:
  • Annotated group labels. Each cluster is defined based on the spurious attribute (gender), resulting in exactly two clusters for CelebA. We sample an equal number of data points from each group $\mathcal{G}=\mathcal{Y}\times\mathcal{A}$ in the training set, creating a group-balanced $\mathcal{D}_{\text{task}}$.
  • Pseudo labels by ERM predictions. Following previous work (Nam et al., 2020; Park et al., 2023; Zhang et al., 2022), we track the ERM predictions on the training set using an early epoch checkpoint during the training process (early-stopped ERM). Samples misclassified by the early-stopped ERM form the minority group, which is assumed to be less influenced by spurious correlations, while the remaining samples form the majority group. We construct $\mathcal{D}_{\text{task}}$ by sampling an equal number of data points from each combination of groups, $\{\text{minority, majority}\}\times\mathcal{Y}$. There are exactly two clusters defined by the minority and majority groups, and $\mathcal{D}_{\text{task}}$ is also constructed class balanced.
Table 10: Details of group accuracy (%) in CelebA dataset when applying PruSC with different defined clustering methods for constructing $\mathcal{D}_{\text{task}}$. Average accuracy of baseline ResNet18 is 89.4%. After pruning, subnetworks has keep-ratio around 50% of ResNet18 weights.
$\mathcal{G}_{1}$ $\mathcal{G}_{2}$ $\mathcal{G}_{3}$ $\mathcal{G}_{4}$
Baselines
ERM 78.5 98.9 98.9 50.2
GroupDRO (fine-tuning on balanced subset) 88.5 91.1 93.5 88.8
PruSC’s contrastive loss with pseudo-labels
Annotated group labels 84.8 90.4 94.8 87.7
ERM predictions 68.7 92.3 91.4 64.9
PruSC 90.5 90.8 89.6 91.2
Table 10 shows the behavior of our subnetwork extraction method with different clustering definitions. The results show that our method can be used in combination with a wide range of clustering methods, all of which improve the worst group accuracy. In particular, the contrastive loss is specifically designed to push apart samples with identical spurious attributes, i.e., samples grouped within the same cluster. Consequently, clusters defined by ERM predictions show limited improvement, as ERM misclassifications tend to divide the dataset into broad minority and majority groups, failing to effectively distinguish between different spurious attributes. On the other hand, clusters defined by ground truth group labels significantly increase the worst-group accuracy (50.2% to 87.7%). Furthermore, the subnetwork achieves similar group performance when applying GroupDRO from scratch.
We hypothesize that the success of PruSC comes from the alignment between the clustering method and the defined contrastive loss. The contrastive loss aims at reshaping the latent space by pulling and pushing samples according to specific criteria. In this context, defining positive and negative samples by representation distances poses the best choice as it directly reinforces the representation structure learned through the contrastive objective. Meanwhile, other sampling approaches that do not use the representation distance information could disrupt this alignment.